OpenStudy (anonymous):

what is the angle btwn the 2 vectos: u=(2,-1) v=(2,0)

OpenStudy (rulnick):

The angle is atan(1/2) = approx 26.565 degrees.

OpenStudy (anonymous):

ohh ill give u the formula we need to use

OpenStudy (anonymous):

cos(x)=(u.v)/(||u||||v||)

OpenStudy (rulnick):

Oh, then you want u . v = 4 + 0 = 4 ||u|| = sqrt(5) ||v|| = sqrt(4) = 2 cos(x) = 4/(2sqrt(5)) x = acos(4/(2sqrt(5)) = 0.46365 radians = 26.565 degrees.